OpenStudy (anonymous):

find the real or imaginary solutions by factoring x^4-10x^2+9=0

OpenStudy (anonymous):

\[x^4-10x^2+9=(x^2-9)(x^2-1)=(x-3)(x+3)(x-1)(x+1)=0\]